A solution is prepared by dissolving 200.0 g of iron (III) chloride (162.2 g/mol) in 200.0 mL of water. What is the molarity of the solution?
Answer #1

Moles of FeCl3 = mass / molar mass = 200 / 162.2 = 1.233 moles

molarity = moles / Volume in liters = 1.233 / 0.200 = 6.16 M